### v1

# Serial and Parallel Two-Column Probing for Mixed-Integer Programming

Abstract
Probing in mixed-integer programming (MIP) is a technique of temporarily fixing variables to discover implications that are useful to branch-and-cut solvers. Such fixing is typically performed one variable at a time—this paper develops instead a two-column probing scheme that instead fixes a pair of variables per iteration. Although the scheme involves more work per iteration compared to the one-column approach, stronger implied bounds as well as more conflicts identified may compensate. Indeed, our prototype implementation was awarded first prize at the MIP Workshop 2024 Computational Competition on novel presolving approaches. This paper presents the aforementioned (serial) prototype and additionally develops an efficient parallelization, leveraging hardware acceleration to further improve overall solve times. Compared to serial two-column probing, our parallel version sacrifices some strength per-pair probed in exchange for greatly increasing the total number of such probings; computational experiments demonstrate its promise.
1 Introduction
For mixed integer programming (MIP) problems, presolving (or preprocessing) is a set of techniques run prior to the main branch-and-bound search that aims to transform a given instance into an equivalent but (hopefully) easier-to-handle formulation. Techniques include eliminating redundancies and strengthening the formulation [6, 24]. Presolve plays an important role in practical results for MIP solvers [8], as well as linear programming (LP) [9, 29], and mixed integer nonlinear programming (MINLP) [42, 39]. Besides general solvers, there are also problem-specific preprocessing methods (see, e.g. [18, 16, 31]).
Probing plays an important role in MIP presolves with demonstrated impact on solvers, e.g. [6]. This technique temporarily fixes binary variables to obtain a variety of useful implications such as variable bound strengthening, and aggregations [45, 7, 2, 48, 32, 26, 1, 6]. The MIP version of probing follows probing for SAT solvers [37], and has subsequently also found use in extensions such as MINLP [35] and multi-stage integer programming [38, 43].
The standard approach to probing involves fixing one binary variable at a time and running e.g. domain propagation [6, Sec. 7.2] and [2, Sec. 10.6]. Probing all binary variables can be potentially time-consuming due to the quadratic runtime complexity associated with the number of nonzero constraint coefficients; thus, solvers typically set a workload-limit and attempt to prioritize variables with a probing order [2, 4, 3, 6, 27]. In this paper, we propose a new two-column probing approach, involving pairs of variable per iteration. As with standard probing, in two-column probing we rely on workload-limits and judicious selection of (pairs of) variables. Our approach is a natural generalization of one-column probing that follows a similar philosophy applied to other parts of MIP such as: multiple columns to reduce models [14, 6]; multiple-row reduction [5]; matrix sparsification [17, 29, 6]; and two-row/column bound strengthening [26].
Additionally, we develop a parallelization scheme for two-column probing that enables substantially more pairs to be processed, but may generate weaker (local) per-probing results compared to the (global) serial scheme. Parallelization has become increasingly important component of MIP solvers [15, 12, 8, 33]; we note also a growing literature in massively parallel computing for MIP (e.g. [41, 23, 46, 34, 47, 40]). Notably, Gleixner, Gottwald, and Hoen [27] parallelized one-column probing by distributing each task of probing a variable to a different thread. In their scheme, global variable bounds, aggregations, and the implication graph must be updated after each iteration; subsequently, the updates involve substantial (time-consuming) communication. For two-column probing, we develop a different approach in Sec. 4 that partitions the set of binary variables and distributes probing work across processors with lower communication costs. We note also that this paper shares the philosophy of our preceding work on parallel conflict graph management [21]: to develop intensive versions of existing preprocessing methods for which the overhead can be mitigated by parallel computing with carefully managed communication costs.
The remainder of the paper is organized as shown in Chart 1. Sec. 2.1 introduces the notation, and Sec. 2.2 presents some simple pre-presolves we used before conducting two-column probing. After simple pre-presolves, there are serial two-column probing shown in Sec.3 and the parallel version in Sec.4. Sec. 5 describes numerical experiments. Sec. 6 concludes.
Sec. 2.1: NotationSec. 2.2: Simple Pre-PresolvesSec. 3: Serial Two-Column ProbingSec. 4: Parallel Two-Column ProbingSec. 5: Numerical ExperimentsSec. 6: Conclusion
Chart 1: Organization of the Paper
2 Setup
2.1 Notation and Basic Definitions
Consider a mixed integer program (MIP) in the following generic form:
$\min$ $c^{T}x$ (1)
s.t. $Ax\circ b$
$\ell\leq x\leq u$
$x_{j}\in\mathbb{Z}\mbox{ for all }j\in\mathcal{I}$
[…]
A coupling matrix $CM\in\mathbb{Z}_{+}^{|\mathcal{B}|\times|\mathcal{B}|}$, where $CM_{i,j}$ is the number of constraints containing binary variables $x_{i}$ and $x_{j}$ with $i,j\in\mathcal{B}$, and $CM_{i,i}=0$. A set packing (clique) constraint is defined as
$\sum_{j\in\mathcal{B}_{S}}x_{j}\leq 1,$
for some $\mathcal{B}_{S}\subseteq\mathcal{B}$. Set packing constraints can, in turn, be inferred from the general MIP formulation. For a given mixed-integer constraint $A_{i\cdot}x\circ_{i}b_{i}$, if $\circ_{i}$ is $\leq$, then we can extract a pure binary variables constraint (PBC) denoted as $\mbox{PBC}(i)$:
$\sum_{j\in\mathcal{B}:A_{ij}>0}A_{ij}x_{j}-\sum_{j\in\mathcal{B}:A_{ij}<0}A_{ij}\bar{x}_{j}\leq b_{i}-\inf\{\sum_{j\not\in\mathcal{B}}A_{ij}x_{j}\}-\sum_{j\in\mathcal{B}:A_{ij}<0}A_{ij}.$ (2)
Note that if $\circ_{i}$ is $\geq$, we can rewrite the constraint as $-A_{i\cdot}x\leq-b_{i}$, and if $\circ_{i}$ is $=$, we can split the constraint into two constraints $A_{i\cdot}x\leq b_{i}$ and $-A_{i\cdot}x\leq-b_{i}$. So we only consider constraints with the form of $A_{i\cdot}x\leq b_{i}$ in the remainder of this section. Let
$\mathcal{C}:=\{i\ |\ i\in\mathcal{M},\mbox{ PBC}(i)\mbox{ is a clique constraint}\},$
which called as clique index set. Let $\mathcal{S}_{sp}:=\{\mbox{PBC}(i)\ |\ i\in\mathcal{C}\}$, which is a collection of all clique constraints, while $\mathcal{S}_{nsp}:=\{A_{i\cdot}x\circ_{i}b_{i}\ |\ i\not\in\mathcal{C}\}$ as a complementary set. A clique table $CT$ includes $|\mathcal{B}|$ sets, where
$CT[j]:=\{i\in\mathcal{M}\ |\ A[i,j]\neq 0,j\in\mathcal{C}\}.$
The conflict number
$conf_{j}=\sum_{i\in CT[j]}\|A[i,:]\|_{0},$ (3)
is the summation of numbers of nonzero elements in constraints $i\in CT[j]$, where $CT[j]$ represents the cliques containing variable $j$. The conflict number is used to estimate the number of conflicts associated with $x_{j}$. The aggregation form $AF$ stores relations like $x_{k}=ax_{j}+b$, and the implication graph $IG$ stores relations like $x_{j}=1\rightarrow x_{k}\geq b$.
2.2 Simple Pre-Presolves
At the preprocessing stage, given a $\mbox{MIP}=(\mathcal{M},\mathcal{N},\mathcal{I},A,b,c,\circ,\ell,u)$, we perform some pre-presolves to detect set packing constraints and constructing $CM$ and $CT$ for subsequent probing, which is described in Alg 1 and Alg 2. This applies well-studied techniques (see e.g. Achterberg et al. [6]).
Input: $\mbox{MIP}=(\mathcal{M},\mathcal{N},\mathcal{I},A,b,c,\circ,\ell,u)$
Output: MIP after applying Simple Presolve, $\mathcal{S}_{sp}$, $\mathcal{S}_{nsp}$, $\mathcal{C}$
1 Set $\mathcal{S}_{sp}$, $\mathcal{S}_{nsp}:=\emptyset$;
2 Remove empty constraints and singletons from MIP and conduct a one-round single-row bound strengthening to MIP;
3 Check integer variable bounds and identify binary variables;
4 for $i\in\mathcal{M}$ do
5       if $\mbox{PBC}(i)$ is a set packing constraint then
6             Set $\mathcal{S}_{sp}:=\mathcal{S}_{sp}\cup\{\mbox{PBC}(i)\}$;
7             Set $\mathcal{C}:=\mathcal{C}\cup\{i\}$;
8      else
9             Set $\mathcal{S}_{nsp}:=\mathcal{S}_{nsp}\cup\{A_{i\cdot}x\leq b_{i}\}$;
10
11       end if
12
13 end for
return MIP after applying Simple Presolve, $\mathcal{S}_{sp}$, $\mathcal{S}_{nsp}$, $\mathcal{C}$.
Algorithm 1 One-round Simple Presolve
All techniques mentioned in line 2 of Alg. 1 can be found in [6, Sec 3.1 and 3.2]. Empty constraints and singletons are discarded as they cannot be probed; moreover, the strengthening of variables bounds allows for detection of binary variables that are otherwise encoded as general integers. The complexity of Alg 1 is $O(NNZ)$, where $NNZ$ is the number of nonzero elements in $A$ [27].
Input: $\mbox{MIP}=(\mathcal{M},\mathcal{N},\mathcal{I},A,b,c,\circ,\ell,u)$, $\mathcal{S}_{sp}$, $\mathcal{S}_{nsp}$, $\mathcal{C}$
Output: $CM$, $CT$
1 Set $CM\in\mathbb{Z}_{+}^{|\mathcal{B}|\times|\mathcal{B}|}$ as all-zero sparse matrix, $CT:=\cup_{j\in\mathcal{B}}\{\emptyset\}$;
2 for $s\in\mathcal{S}_{nsp}$ do
[…]
4             for all pairs of $(x_{i},x_{j})$ from $s$ do
5                   $CM_{i,j}:=CM_{i,j}+1$;
6
7             end for
8
9       end if
10
11 end for
12for $i\in 1:|\mathcal{C}|$ do
13       Set $s,j$ as the $i$-th element of $\mathcal{S}_{sp}$ and $\mathcal{C}$ separately;
14       for $x_{k}\in s$ do
15             Set $CT[k]:=CT[k]\cup\{j\}$;
16
17       end for
18
19 end for
return $CM$, $CT$.
Algorithm 2 Building $CM$ and $CT$
We use a sparse matrix to represent $CM$, which has quadratic memory complexity associated with the length of cliques. To avoid too much memory consumption, we set a size_limit in line 3 of Alg. 2 and exclude cliques larger than the size_limit. We also set a total workload in $CM$ constructing, e.g. the number of nonzero terms processed by our method is at most work_limit. The specific values of these work-limit parameters in our experiment will be provided in Sec. 5.2. In line 10, $s$ is the $i$-th clique in $\mathcal{S}_{sp}$ and $j$ is the $i$-th index in $\mathcal{C}$.
Specifically, because current solvers [30, 2, 25, 20] have already implemented Alg. 1 and it is fast, we excluded its runtime in experiments (which is marginally small regardless). However, we include the runtime of Alg. 2 because different solvers may construct clique table (or conflict graph [10, 1]) with different data structures, which makes various building runtime and may impact subsequent probing runtime. And $CM$ may not be used in other presolve routines.
3 Two-column Probing (serial)
Chart. 2 summarizes our two-column probing method. The remainder of this section provides details on each component.
Input MIPPre-Presolves, see Sec. 2.2, Alg. 1Construct $CM,CT$, see Sec. 2.2, Alg. 2Determine Probing Order, see Sec. 3.1, Alg. 3Two-Column Probing, see Sec. 3.2, Alg. 4Build Reduced Model, see Sec. 3.4Output Reduced MIP
Chart 2: Two-column Probing
3.1 Selecting Variable Pairs
When considering the order of variables to probe, Achterberg [2] determines a probing order of binary variables based on three factors: the number of constraints containing a binary variable $x_{j}$, the number of arcs of $x_{j}$ in the implication graph, and the number of conflicts between $x_{j}$ and other binary variables. For our implementation we focus on constructing an independent probing subroutine, and so we do not generate an implication graph prior to running probing; in practice the implication graph is generated from other preprocessing procedures and refined during probing. Thus, we define probing order in a simpler way, estimating the number of conflicts of binary variable $x_{j}$ by conflict number defined in Equation (3) by $CT$ from Alg. 2. The number of constraints containing $x_{j}$ is $\|A[:,i]\|_{0}$, where the zero-norm represents the number of nonzero elements from one vector. Because pairs of variables are considered, besides the number of constraints containing one binary variable, we consider the number of constraints containing two binary variables. This number is calculated via the coupling matrix $CM\in\mathbb{Z}_{+}^{|\mathcal{B}|\times|\mathcal{B}|}$ from Alg. 2.
Alg. 3 determines probing order. Line 1 selects Cand_Number candidate pairs of variables via binary search. $conflict$ in lines 3-8 acts a penalty to help avoid conflicting variables in the probing order—conflicts reduce the number of useful combinations among the pair of binary variables, i.e. $(0,0),(0,1),$ $(1,0),(1,1)$. Line 10 fetches up to a fixed number of pairs, Max_Probe_Number, for probing, and applies partial sorting (see e.g. [44]).
Input: $CM$, $CT$
Output: Indices of candidate binary variables for the probing method
1 Set $Cand$ as an index set including indices of largest Cand_Number cells from $CM$;
2 for $(i,j)\in Cand$ do
3       if $CT[i]\cup CT[j]\neq\emptyset$ then
4            Set $conflict=10$;
5      else
6             Set $conflict=1$;
7
8       end if
9      Set $score(Cand_{i,j})=\frac{10CM_{ij}+\|A[:,i]\|_{0}+\|A[:,j]\|_{0}+3conf_{i}+3conf_{j}}{conflict}$;
10
11 end for
12Partially sort $Cand$ to extract the top Max_Probe_Number elements with $score(Cand_{i,j}^{(1)})\geq score(Cand_{i,j}^{(2)})\geq...\geq score(Cand_{i,j}^{(1000)})$;
return Top $10^{3}$ elements extracted in Line 5.
Algorithm 3 Sort Variables for Probing
3.2 Two Column Probing
After selecting index pairs, we probe every selected pair of binary variables with Alg. 4.
Input: MIP, candidate pairs of variables $Cand$, $CT$, $ub,lb$
Output: $ub,lb$, $AF$, new conflicts set $New\_Conflict$
1 Set $AF,IG,New\_Conflict:=\emptyset$;
2 for $(x_{i},x_{j})\in Cand$ do
3       If both $x_{i}$ and $x_{j}$ have been probed, go to next iteration;
4       for $v$ = $(0,0),(1,0),(0,1),(1,1)$  do
5             Fix $x_{i},x_{j}=v_{1},v_{2}$;
6             Propagate $CT,AF,IG$ to obtain temporary variable bounds $ub_{temp},lb_{temp}$;
7             Call domain propagation to obtain implied bounds $ub(v),lb(v)$;
8             If the problem is infeasible, add $(2v_{1}-1)x_{i}+(2v_{2}-1)x_{j}\leq v_{1}+v_{2}-1$ to $CT$ and $New\_Conflict$;
9
10       end for
11      If there are two or more values of $v$ are infeasible, fix $x_{i}$ or $x_{j}$;
12       Update $ub=\min\{ub,\max\{ub(v)\}\}$, $lb=\min\{lb,\max\{lb(v)\}\}$;
13       Update $AF,IG$ for $x_{i}$ and $x_{j}$ separately;
14       Update $IG$ for $(x_{i},x_{j})$;
15       Check termination condition;
16
17 end for
Return $ub,lb$, $AF$, $New\_Conflict$.
Algorithm 4 Two-column Probing
In the input, $Cand$ includes Max_Probe_Number pairs of binary variables selected from Alg. 3. Line 7 applies domain propagation (see e.g. [6, Sec. 3.2]) to get implied variable bounds. In line 8, $(2v_{1}-1)x_{i}+(2v_{2}-1)x_{j}\leq v_{1}+v_{2}-1$ cuts off $x_{i}=v_{1},x_{j}=v_{2}$, which is a detected clique. Line 10 fixes variables as follows:
  • Fixes $x_{i}=1$ if both $x_{i}=0,x_{j}=1$ and $x_{i}=0,x_{j}=0$ are infeasible. Likewise with the complement, $x_{i}=0$.
  • Fixes a pair of a variables if the other 3 of 4 assignments are infeasible. For instance, if $(x_{i},x_{j})\in{(0,0),(1,0),(0,1)}$ are infeasible then $x_{i}=1,x_{j}=1$.
  • Replace $x_{j}$ with $x_{i}$ in the formulation if both $x_{i}=0,x_{j}=1$ and $x_{i}=1,x_{j}=0$ are infeasible (i.e. $x_{i}=x_{j}$). Likewise check if this holds for the complement: $x_{i}=1-x_{j}$.
Line 11 updates global variable bounds. In line 12, we update $AF$ and $IG$ for each variable in a pair separately. For a given variable $x_{i}$ we set,
$ub(x_{i}=1):=\max\{ub(1,1),ub(1,0)\},lb(x_{i}=1):=\min\{lb(1,1),lb(1,0)\},$
$ub(x_{i}=0):=\max\{ub(0,1),ub(0,0)\},ub(x_{i}=0):=\min\{lb(0,1),lb(0,0)\}.$
Aggregation and implication are then handled as in the standard one-variable probing [6, Sec. 7.2] or [2, Alg. 10.12, step 4 (c)(d)]. After line 12, we also add implications like $x_{i}=1,x_{j}=1\rightarrow x_{k}\geq b$ to $IG$. To avoid redundancy, pair implications are only added if the individual implications $x_{i}=1\rightarrow x_{k}\geq b$ or $x_{j}=1\rightarrow x_{k}\geq b$ are not in $IG$.
We defer detailed discussion of the termination condition in line 14 to Sec. 3.3.1.
3.3 Discussion and Analysis
In terms of computational complexity, within the for-loop in line 2 of Alg. 4, domain propagation is in $O(NNZ)$, and updating variable bounds, $AF,IG,CT$ are in $O(n)$. Hence, the per-iteration complexity is the same as in one-column probing [27]. The maximum number of iterations is $O(|B|^{2})$ (governed by the number of possible pairs), whereas standard one-column probing is applied once per variable and so has $O(|B|)$ iterations; in either case, workload limits are imposed to make this a constant factor in practice. Note that $\lceil\frac{B}{2}\rceil$ pairs in $Cand$ are sufficient to attain the same (or better) results as one-column probing, provided such pairs do not share variables.
Practically speaking, the most time-consuming step in probing (both single- and two-column) is generating $CT,AF,IG$, and domain propagation. Alg. 4 calls $CT,AF,IG$, and domain propagation $4$ times, which is twice the computational effort of single-column probing, i.e. running single probing to each $x_{i}$ and $x_{j}$ separately. One benefit of this is that all aggregations and implications from single-column probing will also be detected in line 12. The reason is that the implied bounds in line 12 are no weaker than the implied bounds from the one-column probing. For the variable upper bound, suppose $v_{1},v_{2}\in\{0,1\}$,
$ub(x_{i}=v_{1},x_{j}=v_{2})\leq ub(x_{i}=v_{1},x_{j}\in\{0,1\}),$
$ub(x_{i}=v_{1},x_{j}=1-v_{2})\leq ub(x_{i}=v_{1},x_{j}\in\{0,1\}),$
$\implies$ $\max\{ub(x_{i}=v_{1},x_{j}=v_{2}),ub(x_{i}=v_{1},x_{j}=1-v_{2})\}\}$
$\leq$ $ub(x_{i}=v_{1},x_{j}\in\{0,1\});$
thus the implied upper bound from two-column-fixing is no weaker than the bound from one-column-fixing. The lower bound can be shown similarly. We also provide an example to show that two-column probing may return a stronger variable bound in domain propagation. Consider the following system:
$t+x+y+z\geq 1,$ (4)
$w+y\leq 1,$ (5)
$w-z\geq 0,$ (6)
$w,x,y,z\in\{0,1\},t\in[-3,3]$ (7)
Fixing $x=1$ implies $t\geq-2$ from Eq. (4). However, fixing either $x=1,w=1$ or $x=1,w=0$, implies $t\geq-1$, and so $x=1$ implies $t\geq-1$. Additionally, line 8 in Alg. 4 can detect some conflicts between binary variables and line 13 can provide two-variable implications that cannot be obtained by single-column probing.
However, although two-column probing can provide possibly stronger (and never weaker) reduction per-pair compared to one-column probing, not all variables are considered: in line 1 of Alg. 3, only pairs of variables appearing in at least one constraint together are considered. Thus, some binary variables that only appear in constraints alone are ignored by two-column probing; indeed, in experiments we observe that this constitutes nearly half of all binary variables. This is an implementation choice that allows us to easily test the effects of two-column probing above-and-beyond those of one-column probing.
3.3.1 Termination Conditions
[…]
3.4 Output Management
Two-column probing provides strengthened variable bounds, new conflicts, $AF$, and $IG$. We modify the original MIP by replacing the variable bounds with the strengthened bounds and adding clique cuts from new conflicts and aggregation constraints like $x_{i}=ax_{j}+b$ from $AF$. However, the number of implications from $IG$ is quite large, so we discard them; in a more mature implementation there may be a better management scheme that can pass $IG$ in whole or in part in a productive way.
4 Parallelized Two-Column Probing
In this section, we parallelize our two-column probing procedure. We adopt a different approach from that of Gleixner, Gottwald, and Hoen [27] because the two-column probing method involves more global bound updates, aggregations, and implications than one-column probing, and so the communication cost of their scheme may be prohibitive in the extended version. Instead, our method partitions the binary variables and assigns a partition to each thread, which can then be handled independently (Sec. 4.1). Locally, two-column probing (Alg. 4) is applied to variables assigned to each thread. Then the results are reduced/merged globally by using a customized implication analysis procedure (Sec. 4.2).
A diagram of of our parallel probing is given in Chart. 3. Note that, if run over the same pairs of variables, the results will tend to be weaker than the serial two-column method: in order to reduce communication costs, our per-thread probing is conducted locally, which could potentially miss out on more effective global updates (i.e. interactions between variables in different threads). The trade-off between effectiveness per-probing and parallel efficiency (total number of probes conducted) is governed by problem structure.
[…]
Chart 3: Two-column Probing (parallel). Orange color rectangles represent new serial subroutines. Gray colors are run in parallel, with each thread applying the algorithms in the dashed rectangle.
4.1 Variable Partitioning
The main interdependence between the outer for-loop iterations (line 2 of Alg. 4) of the serial implementation occurs in the inner for-loop (line 4). After setting two binary variables to their implied bounds, two-column probing propagates $CT$ to fix additional binary variables. If these $CT$-fixing variables have been probed in previous iterations, two-column probing may fix more variables via $AF$ and $IG$ associated with these $CT$-fixing variables. Thus, for parallelization, we aim to assign variables that conflict with each other to the same thread. Note, however, that maximizing the number of conflicts assigned to each thread is an NP-hard partitioning problem [19, 22]—hence, we propose a greedy heuristic partitioning method described in Alg. 5.
Input: $\mathcal{B}$, $\mathcal{S}_{sp}$, thread number $k$
Output: Partitioned subsets
1 Set $\mathcal{I}_{j}:=\emptyset$ for $j=1,...,k$;
2 Set $assigned_{i}:=0$ for $i\in\mathcal{B}$;
3 Set $j=0$;
4 for $clq\in\mathcal{S}_{sp}$ do
5       for $i$-th variables $\in clq$ do
6             if $assigned_{i}\leq 0$ then
7                   Set $\mathcal{I}_{j+1}:=\mathcal{I}_{j+1}\cup\{i\}$;
8                   Set $assigned_{i}:=1$, $j:=(j+1)\mod{k}$;
9
10             end if
11
12       end for
13
14 end for
15
16for $i\in\mathcal{B}$ do
17       if $assigned_{i}\leq 0$ then
18             Set $j^{*}:=\arg\min\{\|\mathcal{I}_{j}\|,j=1,...,k\}$;
19             Set $\mathcal{I}_{j}:=\mathcal{I}_{j}\cup\{i\}$;
20
21       end if
22
23 end for
return $\mathcal{I}_{j}$ for $j=1,...,k$
Algorithm 5 Assign Variables to Different Threads
Algorithm 5 assigns variables from the same clique to the same thread, as shown in Fig. 1, where a red dashed rectangle indicates binary variables assigned to one thread. If a variable has already been assigned, the heuristic does not reassign it (lines 6 and 13).
Thread 1Thread 2Thread 3123456789
Figure 1: Conflict Graph Constructed from 4 Cliques
4.2 Implication Analysis
In one-column probing, a follow-up preprocessing method called implication graph analysis [2] can use the results of probing to update global bounds and detect new aggregations. Implication analysis is similar to one-column probing but drops domain propagation to save time. However, this procedure is notably less effective for serial two-column probing: based on experiments, we observed that implication graph analysis detects very few new aggregations as our two-column probing tends to probe individual variables multiple times, resulting in diminishing returns. However, our parallel two-column algorithm ignores pairs of variables between two threads, and so we adapt implication analysis by running it only on these neglected pairs to try and recover some of the advantage of global probing. After running local two-column probings, we first update global variable bounds by $ub:=\min_{i=1,...,k}\{ub^{i}\}$ and $lb:=\max_{i=1,...,k}\{lb^{i}\}$, where $ub^{i},lb^{i}$ are the upper and lower bounds from the $i$-th thread. We then conduct implication analysis on pairs of variables between threads as described in Alg. 6.
Input: $CT$, $AF$, $IG$, $ub$, $lb$
Output: updated $ub$, $lb$, $AF$
1 Set $\mathcal{P}:=\{i\ |\ x_{i}\mbox{ has been probed}\}$;
2 for $i\in\mathcal{P}$ do
3       for $v=0,1$ do
4             Propagate $CT$ for $x_{i}=v$ to fix conflicting variables;
5             Propagate $AF,IG$ for $x_{i}=v$ and fixed variables to obtain temporary variable bounds $ub_{temp},lb_{temp}$;
6             Update $ub$, $lb$, and $AF$;
7
8       end for
9
10 end for
return $ub$, $lb$, $AF$
Algorithm 6 Implication Analysis
For each $x_{i},i\in\mathcal{P}$ ($\mathcal{P}$ is defined in line 1 of Alg. 6), after setting it to 00 or $1$, Alg. 6 follows a procedure similar to Alg. 4 but drops domain propagation to save time. In line 4, fixing $x_{i}$ and propagating $CT$ may fix some variables that were assigned to different threads, such as nodes $4$ and $5$ in Fig. 1. Subsequently, propagating $AF$ and $IG$ uses local probing outcomes across threads to detect potential new/stronger implications for $x_{i}$. As a result, probing outcomes from different threads are exchanged and merged into the main thread, restoring some of the global effects that were lost by parallelization.
4.3 Discussion and Analysis
Parallel two-column probing is described as Alg. 7 and Chart. 2.
Input: MIP, thread number $k$
Output: Reduced MIP
1 Conduct simple pre-presolve and construct $CT$, $CM$;
2 Assign variables to different threads: $\mathcal{I}_{1},...,\mathcal{I}_{k}$;
3 for $i\in 1:k$ do parallel
4       Determine probing order;
5       Conduct Two-Column Probing for $\{x_{j},j\in\mathcal{I}_{i}\}$;
6
7 end for
8Reduce probing results by Implication Analysis and build reduced MIP;
return MIP
Algorithm 7 Parallel Two-Column Probing
[…]
5 Numerical Experiments
This section describes our implementation and experiments. All code is publicly available at https://github.com/foreverdyz/TwoColumnProbing.
5.1 Experimental Setup
5.1.1 Test Set
Experiments are conducted on the benchmark set of the MIPLIB 2017 Collection [28], consisting of $240$ instances. We exclude $3$ instances due to the memory-handling issues in JuMP. We remove $47$ instances where no pairs of binary variables appears in the same constraint–hence our method terminates with no result. The results below are run on the remaining $190$ instances.
5.1.2 Software and Hardware
All algorithms are implemented in Julia 1.8.5 [13]. We solve both original MIPs and reduced MIPs (MIPs after applying the two-column probing) SCIP v0.11.14 [4, 3] and an alpha (prototype) version of JuMP 1.28 [36] on an Owens node from the Ohio Supercomputer Center. This node includes 2 sockets, with each socket containing 14-core Intel Xeon E5-2680 CPUs. The node has 128 GB of memory and runs 64-bit Red Hat Enterprise Linux 7.9.0. We use $16$ GB, $16$ cores for the parallel computation, and $1$ core for the SCIP solver.
5.2 Parameter Setting
[…]
5.2.1 Measurements
Pre Time measures time spent in two-column probing. This excludes overhead from the reading time of the input file and the time to build and write the reduced model to file by JuMP, but includes building $CM$, $CT$, selecting variable pairs, two-column probing, and the implication analysis. Pairs # recorded how many pairs of binary variables are probed by (parallel) two-column probing.
Org Time is the time reported by SCIP for solving the original models directly read form the input file. Rd Time is the time reported by SCIP for solving the reduced models generated by two-column probing. Runtime Comp. is defined as
[…]
We also record the primal-dual gap based on the primal and dual bounds reported by SCIP. To get a gap within a fixed range, here we use the primal-dual gap defined by Berthold [11]. This gap is between $[0,100]$: if SCIP finds an optimal solution or certificates the infeasibility, the gap is 00; if the solver does not find any feasible solution but also cannot certificate the infeasibility, the gap is $100$; otherwise, the gap is calculated as
$\texttt{GAP}:=\frac{|\texttt{primal\_bound}-\texttt{dual\_bound}|}{\max\{|\texttt{primal\_bound}|,|\texttt{dual\_bound}|\}}\times 100.$ (9)
5.3 Parallel Performance
We present both the geometric mean of probing runtimes (denoted as Pre Time) with a shift of 1 as well the average number of pairs of probed variables (denoted as Pairs #) in Table 1. This data is further illustrated in Fig. 2, and we fit the Pairs # in different number of threads with a sublinear function $f(k):=35k^{\log_{2}(1.45)}$ where $k$ is the number of threads and $\log_{2}(1.45)\approx 0.54<1$. 16 threads allowed for approximately 4 times more pairs to be considered.
Table 1: Parallel Performance: Pre Time is geometric mean with a shift of 1, and Pairs # is the number of pairs of binary variables processed
Threads # 1 Thread 2 Threads 4 Threads 8 Threads 16 Threads
Pre Time 3.2 3.4 3.2 3.2 3.1
Pairs # 38.0 50.2 74.8 106.6 154.5
111122224444888816161616404040406060606080808080100100100100120120120120140140140140160160160160Number of Threads [$k$]Pairs #Parallel Two-Column ProbingFitting Function: $35k^{\log_{2}(1.45)}$
Figure 2: Pairs # vs. different number of threads.
Table 1 and Fig. 2 indicate that (parallel) two-column probing takes $3.1\sim 3.4$ seconds (with a shift of 1), while the number of pairs of probed variables increases with a sublinear rate.
5.4 Default SCIP
[…]
For column labeled “serial”, we report for serial two-column probing: in the first row, the geometric mean of Runtime Comp. (defined in Equation (8)) with a shift of 10 for instances that can be solved by SCIP within the 3600-second time limit on either original or reduced model from serial two-column probing; in the second row, the geometric mean of GAP (defined in Equation (9)) with a shift of 1 for instances that cannot be solved by SCIP within the time limit on either original or reduced model from serial two-column probing. Likewise, for $2,4,8,16$ threads, we report shifted geometric means of Runtime Comp. and GAP.
Table 2: Two-Column Probing + Org vs. Org; Runtime Comp. for within time-limit cases and GAP for exceeding time-limit cases
Threads # Serial 2 Threads 4 Threads 8 Threads 16 Threads
Runtime Comp. 97.6% 100.0% 98.3% 97.5% 95.4%
GAP 88.4% 94.2% 96.5% 93.8% 98.2%
From Table. 2, serial two-column probing attains a $2.4\%$ speedup compared to SCIP with default settings, as well as substantially improving duality gap on unsolved instances. The parallel method has weaker performance on 2 threads, but provides substantial speedups with 16 threads.
In Table. 3 we provide runtime breakdowns by problem difficulty, with the brackets set using default SCIP runtimes. Two-column probing appears more advantageous on harder problems.
Table 3: Two-Column Probing + Org vs. Org; Runtime Comp. for different Org Time brackets.
Brackets (s) # of Instances Serial 2 Threads 4 Threads 8 Threads 16 Threads
All 190 97.6% 100.0% 98.3% 97.5% 95.4%
$\geq 10$ 183 96.4% 98.7% 95.9% 96.1% 93.1%
$\geq 100$ 157 89.9% 89.3% 88.3% 89.3% 86.3%
$\geq 1000$ 125 69.2% 67.2% 71.8% 68.5% 63.3%
5.5 Two-Column vs One-Column Probing
We also compare two-column probing with one-column probing disabled (setting parameter “propagating/probing/maxuseles” as “false”). Namely experiments from Sec. 5.4 were rerun comparing two-column probing + default SCIP without one-column probing vs. default SCIP. Results are summarized in Table 4.
Table 4: Two-Column Probing + Org - One-Column Probing vs. Org; Runtime Comp. for within time-limit cases and GAP for exceeding time-limit cases
Threads # Serial 2 Threads 4 Threads 8 Threads 16 Threads
Runtime 103.1% 99.4% 102.3% 103.0% 103.8%
GAP 109.1% 110.6% 108.1% 114.8% 107.3%
Table. 4 suggests that standalone two-column probing is less effective than one-column probing as implemented in SCIP. Two contributing factors are suggested here. First, two-column probing affects fewer variables, as binary variables that only appear in constraints alone are not probed. We do this to reduce overlapping efforts with one-column probing. Indeed, we have calculated that approximately half of variables are ignored in these MIPLIB instances. Second, our two-column implementation is not as well-integrated as SCIP’s native one-column method. Unlike our code, SCIP’s probing accesses $CT,IG$ and a reduced model that is obtained from other presolve subroutines, and thus probing can be more judiciously be deployed (and with substantially less overhead) by avoiding instances that are more likely to be easy. This is supported by solve time breakouts given in Table 5.
Table 5: Two-Column Probing + Org - One-Column Probing vs. Org; Runtime Comp. for different Org Time brackets.
Brackets Instances # 1 Thread 2 Threads 4 Threads 8 Threads 16 Threads
All 190 103.1% 99.4% 102.3% 103.0% 103.8%
$\geq 10$ 183 104.0% 98.1% 101.4% 103.3% 104.0%
$\geq 100$ 160 96.5% 93.6% 92.5% 96.4% 98.2%
$\geq 1000$ 127 77.0% 66.2% 78.2% 80.9% 80.5%
6 Conclusion
This paper develops the first two-column probing procedure, with both a serial and parallel version provided. Although our basic implementation does not have the maturity of fully-integrated one-column probing, empirical results with SCIP nonetheless indicate the promise of our approach, especially on more difficult MIPLIB instances. The parallel version appears advantageous compared with our serial method using 8 or 16 threads.
For future work, it is worth considering mixed-integer nonlinear problems, where the cost of solving a relaxation tends to be a relatively more expensive compared to the linear setting. In this setting our more intensive version of probing could prove advantageous.